# Causal Classification of Pathological Misner-Type Spacetimes

### v2

@@ Abstract.
We analyze the Misner space, incorporating Kip Thorne’s ‘moving wall’ model, and pseudo-Schwarzschild spacetime, exploring their geometric structures, geodesics, and causal properties. We propose that pseudo-Schwarzschild spacetime serves as a non-flat generalization of Misner space. We then introduce the pseudo-Reissner-Nordström spacetime, a novel chronology-violating model that extends the previous two frameworks and unifies the geodesic structures of all three spacetimes. We suggest that these spacetimes constitute a causality-violating family, interconnected through common fundamental causal features—geodesics, Cauchy and chronology horizons, and acausal regions—leading to a conjecture that they are isocausal.
@@ 1. Introduction
General relativity defines a class of cosmological models, each representing an idealization of a physically possible universe compatible with the theory. A cosmological model is a mathematical description of this idealized universe, referred to as a spacetime and represented by a pair $(M,g)$. Here, $M$ is a differentiable manifold that captures the topology and continuity of the universe, while $g$ is a smooth, symmetric, non-degenerate $(0,2)$-tensor field on $M$. The metric tensor $g$ encodes the geometric and causal structure of spacetime, with each point in $M$ representing a possible event. In this context, the theoretical possibility of time travel—through closed timelike curves that allow a timelike observer to return to an event in his own past—has been studied extensively within general relativity [1, 4, 9, 19, 12, 21, 22]. In 1967 Charles Misner introduced the Misner space [14] as a basic spacetime with a chronology violating region. However, while it is relatively simple to construct spacetimes containing closed timelike curves, finding examples that are also realistic, to some extent, is more challenging. The more sophisticated causality violating pseudo-Schwarzschild spacetime, proposed by Amos Ori [16] in 2007, offers a more realistic alternative to idealized black hole models by incorporating internal features such as Cauchy horizons, avoiding singularities, and reflecting the breakdown of determinism inside black holes. [^1] An "idealized black hole model" is a simplified theoretical depiction of a black hole, commonly assuming perfect spherical symmetry and no rotation, and typically derived from the Schwarzschild metric.
@@
We begin by examining those two seemingly different spacetimes, the Misner space, also incorporating Kip Thorne’s ‘moving wall’ model, and pseudo-Schwarzschild spacetime. We aim to shed light on their geometric structures and causal properties, with a particular focus on pathologies. These two spacetimes are related through their chronological and global properties. Our findings suggest that the pseudo-Schwarzschild spacetime can be viewed as a non-flat generalization of Misner space. By examining the relationship between the pseudo-Schwarzschild and Misner spacetimes, we put forward a conjecture in 2016 that these spacetimes are isocausal [18].
We introduce a novel chronology-violating spacetime, referred to as the pseudo-Reissner-Nordström spacetime. This model represents a further generalization of the previous two spacetime frameworks. We derive the pseudo-Reissner-Nordström spacetime from the well-known Reissner-Nordström spacetime and review our main results within this more general framework. By situating our findings within this broader context, we conclude that these spacetimes—all of which share a similar geodesic structure, along with Cauchy and chronology horizons, and acausal regions—form a causality-violating family.
Ultimately, all of these spacetimes can be expressed in a warped product form, with their metrics decomposing into cylindrical and hyperbolic components. From a purely mathematical standpoint, the resulting similarities in causal properties might seem unsurprising. However, from a physics perspective, these interrelationships are noteworthy. The spacetimes in question differ significantly in their physical interpretation and geometric origin—ranging from a flat vacuum solution of the Einstein equations, to an asymptotically flat black hole solution of the vacuum field equations, to a spacetime that violates the weak energy condition and requires so-called ‘exotic matter’ as its source. Despite these differences, all exhibit analogous causal behavior. This distinction is crucial, as it sets the pseudo-Reissner-Nordström spacetime apart from both the Misner and pseudo-Schwarzschild spacetimes, which are vacuum solutions. The non-vacuum nature of the pseudo-Reissner-Nordström spacetime is precisely what makes its inclusion in the family of causality-violating spacetimes both surprising and significant.
The intricate interrelationships of those spacetimes offer a unified framework for understanding their causal and geometric features, motivating the conjecture that all three spacetimes are isocausal. While the conjecture is supported by our preliminary analysis and the foundational basis outlined in this article, the formal proof will be presented in future work.
@@ 2. Misner space
Misner space [14], typically presented as a $2$-dimensional toy model, is a type of spacetime in general relativity that that illustrates how singularities might form. This spacetime begins with causally well-behaved initial conditions but later develops closed timelike curves, leading to a chronology-violating region. Characterized by a conformally flat metric, Misner space is also Ricci-flat, making it a valid vacuum solution to Einstein’s field equations in the absence of matter and energy. The $4$-dimensional version can be obtained as a straightforward extension. It can be considered as Minkowski space with an altered topology due to identification under a boost. [^3] A Lorentz boost is a Lorentz transformation with no rotation. To see this, start with the Lorentz-Minkowski spacetime $\mathcal{\mathscr{M}}:=(\mathbb{R}^{4},\tilde{\eta})$, defined as the smooth manifold $\mathbb{R}^{1,3}:=(\mathbb{R}^{4},\tilde{\eta})$ equipped with Lorentzian coordinates $(\tilde{t},\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{3})$. These coordinates yield the metric tensor $\tilde{\eta}$, given by
@@
(2.2) $\tilde{t}=2\sqrt{-\eta}\cosh(\frac{X_{1}}{2}),\;\tilde{x}_{1}=2\sqrt{-\eta}\sinh(\frac{X_{1}}{2}),\;\tilde{x}_{2}=X_{2},\;\tilde{x}_{3}=X_{3}.$
This coordinate change results in the Misner metric $ds^{2}=-\eta^{-1}d\eta^{2}+\eta(dX_{1})^{2}+(dX_{2})^{2}+(dX_{3})^{2}$, where $0<\eta<\infty$, $0\leq X_{1}\leq 2\pi$, and $0\leq X_{2}$, $X_{3}\leq\infty$. Due to the periodic nature of the coordinate $X_{1}$, the Misner spacetime is defined on $S^{1}\times\mathbb{R}^{3}$. Accordingly, Minkowski space $\mathscr{M}$ serves as the associated covering space of Misner space, where identical points are determined by the identification
@@
with $a$ is the period associated with the periodic direction $X_{1}$. [^4] Note, that these points correspond to $(\eta,X_{1},X_{2},X_{3})\leftrightarrow(\eta,X_{1}+na,X_{2},X_{3})$ in Misner coordinates. However, at $\eta=0$ the metric exhibits an apparent (coordinate) singularity. To address this, we introduce a new set of coordinates to extend the metric beyond $\eta=0$. Define $\varphi=X_{1}-\ln(\eta)$, which transforms the metric into
(2.4) $ds^{2}=2dTd\varphi+Td\varphi{}^{2}+(dX_{2})^{2}+(dX_{3})^{2},$
where $T$ is a time-like coordinate and $\varphi$ is an angular coordinate, with domains $-\infty<T<\infty$ and $0\leq\varphi\leq a$ [^5] Due to the periodicity condition, $\varphi$ can take any real number value. Specifically, $\varphi$ and $\varphi+na$ are equivalent for any integer $n$.
Note that this is only one of two possible inequivalent extensions of Misner space. Alternatively, we could start with the Minkowski metric (2.1) again, but apply a different coordinate transformation, subsequently extending the resulting Misner space across the coordinate singularity in a different way. In both cases, we obtain a cylindrical Misner spacetime defined on $M=S^{1}\times\mathbb{R}^{3}$.
We time orient the Misner space by requiring $-\partial_{T}$ to be future-pointing. It is evident that $\partial_{T}$ is always lightlike, while $\partial_{\varphi}$ is spacelike for $T>0$, lightlike for $T=0$, and timelike for $T<0$, respectively. At $T=0$, there is a null surface, called the chronology horizon, which separates the region of chronology violation from the chronal region. Every causal curve intersects this surface exactly once, making it the Cauchy horizon for any initial hypersurface with $T_{0}=\textrm{const}>0$. Any hypersurface where $T=\textrm{const}>0$ is spacelike and can be chosen as an initial hypersurface for specifying initial data.
@@ 2.1. Kip Thorne’s physics approach to Misner space
Figure 1. Depicted are two spacetime diagrams for Misner space, illustrating the occurrence of closed timelike curves (CTCs). The symbol $\tau$ denotes proper time, and it is important to note that identical values of proper time on both axes are physically identical. CTCs appear due to the time dilation effects of special relativity. The light cone structure is shown in Figure (b). Through every event to the future of the chronology horizon, there exist CTCs, whereas there are no CTCs to the past of the chronology horizon.
Consider the flat $2$-dimensional Minkowski spacetime with a reference frame in Lorentzian coordinates $(t,x)$, where $t$ is interpreted as the time direction and $x$ as the space direction. Next, we identify the $t$-axis at $x=0$ with the $t$-axis at $x=P$ (which we will henceforth refer to as the $\bar{t}$-axis): $(t,P)\sim(t,0)$. At $t=0$, we set the $\bar{t}$-axis into motion with constant speed $\beta$ towards the $t$-axis, while the reference frame in Lorentzian coordinates remains at rest with respect to the $t$-axis. We can think of the $t$-axis as representing an observer at rest and the $\bar{t}$-axis as representing an observer moving at speed $-\beta$ relative the reference frame. This moving observer is equipped with a clock, and $\tau$ denotes the proper time measured by this clock. [^6] Proper time, also called clock time or process time, is a measure of the amount of physical process that a system undergoes. This $x,t$-coordinate system is furnished with the flat Minkowski metric $\widetilde{\eta}$, and therefore the light cones do not tilt or open. Due to the special relativistic time dilation, a region with closed timelike curves develops, see Figure 1.
Note that the worldline of the clock at rest (along the $t$-axis) and the worldline of the clock in motion (along the $\bar{t}$-axis) are identified by $\sim$, meaning the events along these lines are physically identical. However, due to special relativity, the observer at rest perceives the moving clock to tick more slowly. On the $t$-axis we have $\tau=t$, while on the $\bar{t}$-axis, due to time dilation, $\tau=\frac{t}{\sqrt{1-\beta^{2}}}=\gamma t$, where $\gamma:=\frac{1}{\sqrt{1-\frac{v^{2}}{c^{2}}}}=\frac{1}{\sqrt{1-\beta^{2}}}$.
@@
where $\xi=\sqrt{\left(\frac{1-\beta}{1+\beta}\right)}$ is the Doppler blueshift. In the covering space (Figure 3) we get the spatial distance between $W_{0}$ and the $\tilde{t}$-axis by the converging series
@@
[…]
@@
To summarize, this ensures that the intersection points $s_{n}$ are not physical singularities, but merely coordinate singularities. This reasoning is supported by the fact that the $2$-dimensional Misner universe can be viewed as a simplified $2D$ wormhole spacetime [12]; both spacetimes share the same covering space. By projecting a wormhole whose mouths move past each other at speed $\beta$ [21, Figure 3c] onto the $x-t$ plane, we obtain a situation similar to that of Misner space, as described above. However, this wormhole spacetime diagram reveals a new feature: a past chronology horizon, $\mathcal{H}_{-}$. Therefore, the chronology-violating region is confined by two chronology horizons. Like the future chronology horizon $\mathcal{H}_{+}$, the past chronology horizon $\mathcal{H}_{-}$ is a null surface.
@@ 2.2. Geodesics for warped Misner product space
The $4$-dimensional Misner space (2.4) can be expressed as a warped product space of the form $M=M_{Z}\times_{r}H^{2}$, where $M_{Z}$ is the $2$-dimensional Misner-type spacetime with coordinates $(T,\varphi)$, and $H^{2}$ is the $2$-dimensional hyperbolic plane, described by coordinates $(\theta,\phi)$. The general form of the metric combines the $2$-dimensional Misner-type geometry, which encodes the non-trivial topological and causal structure, with a $2$-hyperbolic plane that introduces hyperbolic symmetry. The metric can be written as
@@
where $f(T)$ describes the time-dependent geometry of the Misner-type spacetime and $r$ is the warping function. To interpret this, we consider $\mathbb{R}^{2}$ as being represented topologically by the hyperbolic plane $H^{2}$. The $4$-dimensional Misner space can then be described as a warped product $M:=M_{Z}\times_{r}H^{2}$, with the warping function $r$ being a scalar function $r:M_{Z}\rightarrow\mathbb{R}$, depending on the coordinates of the base manifold $M_{Z}$. The Misner space $M_{Z}=S^{1}\times\mathbb{R}$ is equipped with the cylindrical metric $ds^{2}=2dTd\varphi+Td\varphi{}^{2}$. The fiber $H^{2}$, with coordinates $(\theta,\phi)$, has the hyperbolic metric $g_{H^{2}}=d\theta^{2}+\sinh^{2}\theta d\phi^{2}$, scaled by $r^{2}$. Therefore, the metric on the total space $M$ can be written as $ds^{2}=g_{Z}+r^{2}g_{H^{2}}$, which explicitly expands to
@@
The surface $H_{r}^{2}$, also known as a pseudo-sphere, has constant negative Gaussian curvature of $-1$, which provides the desired high symmetry. The upper sheet of the pseudo-sphere is given by the set
@@
Based on the warped geometry of Misner space, we can focus on the embedded cylinders defined by $\theta=\textrm{const}$ and $\varphi=\textrm{const}$. By doing so, we restrict our attention to the $2$-dimensional Misner cylinder $M_{Z}$, which captures the essential features of the Misner-type geometry.
@@
Given the $2$-dimensional Misner metric $ds^{2}=Td\varphi{}^{2}+2dTd\varphi$, null geodesics are described by
@@
[…]

### v3

@@ Abstract.
We investigate three causality-violating spacetimes: Misner space (including Kip Thorne’s “moving wall” model), the pseudo-Schwarzschild spacetime, and a new model introduced here, the pseudo-Reissner-Nordström spacetime. Despite their different physical origins—ranging from a flat vacuum solution to a black-hole-type vacuum solution to a non-vacuum model requiring exotic matter—all three share a common warped-product structure, 22–dimensional cylindrical base metrics of Eddington-Finkelstein type, and fundamental causal features such as Cauchy and chronology horizons, acausal regions, and analogous geodesic behaviour. Building on a conjecture first proposed in 2016, we present a formal proof that the three models are pairwise isocausal on their universal covers and on suitable causally regular regions of their compactified forms. The proof is constructive, providing explicit causal bijections between the models and identifying a concrete deck-equivariance criterion for when the isocausality descends to the compactified spacetimes. In the absence of equivariance, the covering result does not imply global isocausality, and at most a one-way causal relation holds between the compactified models. These results supply a rigorous causal classification linking these spacetimes, placing them within a unified Misner-type family and providing a framework for extending the classification to other causality-violating solutions.
@@ 1. Introduction
General relativity defines a class of cosmological models, each representing an idealization of a physically possible universe compatible with the theory. A cosmological model is a mathematical description of this idealized universe, referred to as a spacetime and represented by a pair $(M,g)$, where $M$ is a differentiable manifold capturing the topology and continuity of the universe, and $g$ is a smooth, symmetric, non-degenerate (0,2)(0,2)–tensor field encoding the geometric and causal structure. Each point in $M$ represents an event, and the causal structure determined by $g$ governs the possible trajectories of particles and light rays. Within this framework, the theoretical possibility of time travel—through closed timelike curves (CTCs) that allow a timelike observer to return to an event in their own past—has been studied extensively [1, 4, 9, 19, 12, 21, 22]. In 1967, Charles Misner introduced the Misner space [14] as a minimal example of a spacetime with a chronology-violating region. While constructing spacetimes containing CTCs is relatively straightforward, finding examples with a plausible physical interpretation is more challenging. The pseudo-Schwarzschild spacetime, proposed by Ori [16] in 2007, is a more sophisticated causality-violating model that incorporates features such as Cauchy horizons, avoids singularities, and models the breakdown of determinism inside black holes. [^1] An “idealized black hole model” is a simplified theoretical depiction of a black hole, commonly assuming perfect spherical symmetry and no rotation, and typically derived from the Schwarzschild metric.
@@
We begin by examining these two apparently different spacetimes—Misner space (also in the form of Kip Thorne’s “moving wall” model) and the pseudo-Schwarzschild spacetime—to clarify their geometric structures and causal properties, with a particular focus on pathologies. The two are linked by their chronological and global features. Indeed, the pseudo-Schwarzschild spacetime can be viewed as a non-flat generalization of Misner space: in the limit as the mass parameter $M\to 0$, the pseudo–Schwarzschild metric reduces to the Misner metric. In earlier work [18] we conjectured that these spacetimes are isocausal in the sense of García-Parrado and Senovilla.
We then introduce a new chronology-violating spacetime: the pseudo-Reissner-Nordström spacetime, which generalizes the previous two by including an electric charge parameter in analogy with the classical Reissner-Nordström solution. Like its predecessors, this model possesses Cauchy and chronology horizons, acausal regions, and a warped-product structure with a 22–dimensional cylindrical base and hyperbolic spatial fibers. Its inclusion is especially noteworthy because it is a non-vacuum solution, in contrast to the vacuum nature of Misner and pseudo-Schwarzschild spacetimes, yet it still exhibits the same characteristic causal pathologies.
A central contribution of this paper is to replace the earlier conjecture with a formal proof that all three spacetimes—Misner, pseudo-Schwarzschild, and pseudo-Reissner-Nordström—are pairwise isocausal on their universal covers, and on suitable open, causally regular regions of the compactified spacetimes. Proposition 7.1 provides explicit causal bijections between the models in Eddington-Finkelstein-type coordinates, and identifies a concrete deck-equivariance criterion under which this isocausality descends to the compactified spacetimes themselves. Corollary 1 and Remark 7.2 clarify that, in the generic case where this criterion fails, one obtains only a one-way causal relation globally, despite the mutual isocausality of the covering spaces. This sharpens the classification of these causality-violating models and establishes, for the first time, a rigorous causal equivalence framework connecting them.
From a mathematical perspective, the shared warped-product structure and conformal relation of the base metrics make these isocausality results natural; from a physical perspective, the fact that such structurally and interpretatively different spacetimes exhibit identical causal behaviour is striking. This work provides a unified geometric and causal framework for understanding these examples, and paves the way for extending the classification to other models—such as the pseudo-Kerr spacetime—that may also belong to this Misner-type, causality-violating family.
@@ 2. Misner space
[…]
@@
(2.2) $\tilde{t}=2\sqrt{\eta}\cosh(\frac{X_{1}}{2}),\;\tilde{x}_{1}=2\sqrt{\eta}\sinh(\frac{X_{1}}{2}),\;\tilde{x}_{2}=X_{2},\;\tilde{x}_{3}=X_{3}.$
This coordinate change results in the Misner metric $ds^{2}=-\eta^{-1}d\eta^{2}+\eta(dX_{1})^{2}+(dX_{2})^{2}+(dX_{3})^{2}$, where $0<\eta<\infty$, $0\leq X_{1}\leq 2\pi$ as an angular coordinate, and $\infty\leq X_{2}$, $X_{3}\leq\infty$. Due to the periodic nature of the coordinate $X_{1}$, the underlying topology is given by $S^{1}\times\mathbb{R}^{3}$. However, at $\eta=0$ the metric exhibits an apparent (coordinate) singularity. To address this, we introduce a new set of coordinates to extend the metric beyond $\eta=0$. Specifically, we define a new coordinate $\varphi=X_{1}-\ln(\eta)$. To transform the metric accordingly, we also require $d\varphi=dX_{1}+\frac{d\eta}{\eta}$. Thus, the metric transforms into $ds^{2}=2d\eta d\varphi+\eta d\varphi^{2}+(dX_{2})^{2}+(dX_{3})^{2}$. Finally, since $\eta$ is a timelike coordinate, we introduce the substitution $T=\eta$, leading to the metric
(2.3) $ds^{2}=2dTd\varphi+Td\varphi{}^{2}+(dX_{2})^{2}+(dX_{3})^{2},$
where $T$ is a timelike coordinate and $\varphi$ is an angular coordinate, with domains $\infty<T<\infty$ and $0\leq\varphi\leq a$. [^4] Due to the periodicity condition, $\varphi$ can take any real number value. Specifically, $\varphi$ and $\varphi+na$ are equivalent for any integer $n$. Note that this is only one of two possible inequivalent extensions of Misner space. Alternatively, we could start with the Minkowski metric (2.1) again, but apply a different coordinate transformation, subsequently extending the resulting Misner space across the coordinate singularity in a different way. In both cases, we obtain a cylindrical Misner spacetime defined on $M=S^{1}\times\mathbb{R}^{3}$.
Accordingly, Minkowski space $\mathscr{M}$ serves as the associated covering space of Misner space, where identical points are determined by the identification
@@
where $a$ is the period associated with the periodic direction $X_{1}$. [^5] Note, that these points correspond to $(\eta,X_{1},X_{2},X_{3})\leftrightarrow(\eta,X_{1}+na,X_{2},X_{3})$ in Misner coordinates.
We orient Misner space in time by requiring $\partial_{T}$ to point toward the future. The vector field $\partial_{T}$ iis everywhere lightlike, while $\partial_{\varphi}$ is spacelike for $T>0$, lightlike at $T=0$, and timelike for $T<0$. At $T=0$, there exists a null surface known as the chronology horizon, which separates the chronology-violating region from the chronal region. This surface is intersected exactly once by every causal curve, thereby serving as the Cauchy horizon for any initial hypersurface with $T_{0}=\textrm{const}>0$. Hypersurfaces of constant $T=\textrm{const}>0$ are spacelike and and can be chosen as initial data hypersurfaces for specifying the evolution of fields.
@@ 2.1. Kip Thorne’s physics approach to Misner space
Figure 1. Depicted are two spacetime diagrams for Misner space, which illustrate the violation of causality. The symbol $\tau$ denotes proper time, and it is important to note that identical values of proper time on both axes are physically identical. CTCs appear due to the time dilation effects of special relativity. The light cone structure is shown in Figure (b). Through every event to the future of the chronology horizon, there exist CTCs, whereas there are no CTCs to the past of the chronology horizon.
Consider the flat 22–dimensional Minkowski spacetime with a reference frame in Lorentzian coordinates $(t,x)$, where $t$ is interpreted as the time direction and $x$ as the space direction. Next, we identify the tt–axis at $x=0$ with the tt–axis at $x=P$ (which we will henceforth refer to as the $\bar{t}$–axis): $(t,P)\sim(t,0)$. At $t=0$, we set the $\bar{t}$–axis into motion with constant speed $\beta$ towards the tt–axis, while the reference frame in Lorentzian coordinates remains at rest with respect to the tt–axis. We can think of the tt–axis as representing an observer at rest and the $\bar{t}$–axis as representing an observer moving at speed $\beta$ relative the reference frame. This moving observer is equipped with a clock, and $\tau$ denotes the proper time measured by this clock. [^6] Proper time, also called clock time or process time, is a measure of the amount of physical process that a system undergoes. This x,tx,t–coordinate system is furnished with the flat Minkowski metric $\widetilde{\eta}$, and therefore the light cones do not tilt or open. Due to the special relativistic time dilation, a region with closed timelike curves develops, see Figure 1.
Note that the worldline of the clock at rest (along the tt–axis) and the worldline of the clock in motion (along the $\bar{t}$–axis) are identified by $\sim$, meaning the events along these lines are physically identical. However, due to special relativity, the observer at rest perceives the moving clock to tick more slowly. On the tt–axis we have $\tau=t$, while on the $\bar{t}$–axis, due to time dilation, $\tau=\frac{t}{\sqrt{1-\beta^{2}}}=\gamma t$, where $\gamma:=\frac{1}{\sqrt{1-\frac{v^{2}}{c^{2}}}}=\frac{1}{\sqrt{1-\beta^{2}}}$.
@@
where $\xi=\sqrt{\left(\frac{1-\beta}{1+\beta}\right)}$ is the Doppler blueshift. In the covering space (Figure 3) we get the spatial distance between $W_{0}$ and the $\tilde{t}$–axis by the converging series
@@
for $\mid\xi\mid<1$. Thus, the left wall of copy 0copy\,0, denoted $W_{0},$ is located at $\tilde{x}=\frac{-D}{1-\xi}$, and the right wall, $W_{1}$, of copy 0copy\,0 is located at
@@
To summarize, this ensures that the intersection points $s_{n}$ are not physical singularities, but merely coordinate singularities. This reasoning is supported by the fact that the 22–dimensional Misner universe can be viewed as a simplified $2D$ wormhole spacetime [12]; both spacetimes share the same covering space. By projecting a wormhole whose mouths move past each other at speed $\beta$ [21, Figure 3c] onto the xtx--t plane, we obtain a situation similar to that of Misner space, as described above. However, this wormhole spacetime diagram reveals a new feature: a past chronology horizon, $\mathcal{H}_{-}$. Therefore, the chronology-violating region is confined by two chronology horizons. Like the future chronology horizon $\mathcal{H}_{+}$, the past chronology horizon $\mathcal{H}_{-}$ is a null surface.
@@ 2.2. Geodesics for warped Misner product space
The 44–dimensional Misner space (2.3) can be expressed as a warped product space of the form $M=M_{Z}\times_{r}H^{2}$, where $M_{Z}$ is the 22–dimensional Misner-type spacetime with coordinates $(T,\varphi)$, and $H^{2}$ is the 22–dimensional hyperbolic plane, described by coordinates $(\theta,\phi)$. The general form of the metric combines the 22–dimensional Misner-type geometry, which encodes the non-trivial topological and causal structure, with a 22–hyperbolic plane that introduces hyperbolic symmetry. The metric can be written as
@@
where $f(T)$ describes the time-dependent geometry of the Misner-type spacetime and $r$ is the warping function. To interpret this, we consider $\mathbb{R}^{2}$ as being represented topologically by the hyperbolic plane $H^{2}$. The 44–dimensional Misner space can then be described as a warped product $M:=M_{Z}\times_{r}H^{2}$, with the warping function $r$ being a scalar function $r:M_{Z}\rightarrow\mathbb{R}$, depending on the coordinates of the base manifold $M_{Z}$. The Misner space $M_{Z}=S^{1}\times\mathbb{R}$ is equipped with the cylindrical metric $ds^{2}=2dTd\varphi+Td\varphi{}^{2}$. The fiber $H^{2}$, with coordinates $(\theta,\phi)$, has the hyperbolic metric $g_{H^{2}}=d\theta^{2}+\sinh^{2}\theta d\phi^{2}$, scaled by $r^{2}$. Therefore, the metric on the total space $M$ can be written as $ds^{2}=g_{Z}+r^{2}g_{H^{2}}$, which explicitly expands to
@@
The surface $H_{r}^{2}$, also known as a pseudo-sphere, has constant negative Gaussian curvature of 1-1, which provides the desired high symmetry. The upper sheet of the pseudo-sphere is given by the set
@@
Based on the warped geometry of Misner space, we can focus on the embedded cylinders defined by $\theta=\textrm{const}$ and $\varphi=\textrm{const}$. By doing so, we restrict our attention to the 22–dimensional Misner cylinder $M_{Z}$, which captures the essential features of the Misner-type geometry.
@@
Given the 22–dimensional Misner metric $ds^{2}=Td\varphi{}^{2}+2dTd\varphi$, null geodesics are described by
@@
$\psi(T)=\frac{\frac{1}{T}\left(\xi-\omega\sqrt{\xi^{2}+T}\right)}{\omega\sqrt{\xi^{2}+T}}dT=\int\frac{1}{T}\left(\frac{\omega}{\sqrt{1+\frac{T}{\xi^{2}}}}-1\right)dT,$